OpenStudy (anonymous):

Help please, medal will be given Linear Equation in One Variable Translate the Word Problem into a System of Equations On a certain day a snack bar sold 120 ham & chicken sandwiches for 740$. of each chicken sandwich was sold at 5$ and each ham sandwiches at 7$. how many of each did they sell? Mina is 2/3 as old as fe, in 20 yrs Mina will be 7/8 as old as fe how old to each?

OpenStudy (anonymous):

For the first one yu'd set up the equation like this: 120=h+c 740=7h+5c Do yu understand how I got that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well basically, yu have to find one before yu find the other. Does it give yu an answer choice?

OpenStudy (anonymous):

nope just that

OpenStudy (anonymous):

Okay Since im not the BEST at describing things, Ill show yu how to get the answer nd yu tell me if yu understand okay?

OpenStudy (anonymous):

ok thanks.

OpenStudy (anonymous):

120=h+c 740=7h+5c The Last One Represents The Total And How Much Each Sandwich cost. ($7 for Ham and $5 for chicken. 120-c=h+c-c So H=120-c Then Plug that into the other problem. 740$=7(120-c)+5c Do yu understand what Im doing?

OpenStudy (anonymous):

ok go on.

OpenStudy (anonymous):

Distributive property is next. 740=(7*120 - 7*c)+5c 740=840-7c+5c 740=840-2c -840 -840 -100=-2c Can yu finish it from there?

OpenStudy (anonymous):

where did the 5c go? 740=(7*120 - 7*c)+5c 740=840-7c+5c 740=840-2c -840 -840 -100=-2c

OpenStudy (anonymous):

Combine The Like Terms. -7c+5c= -2c

OpenStudy (anonymous):

why didnt we transpose 840 to the left?

OpenStudy (anonymous):

I Did. But to get rid of it on one side I had to subtract it. And yu have to do the same thing to both sides.

OpenStudy (anonymous):

Thats how I got the -100

OpenStudy (anonymous):

ok i get it, but how about the last step?

OpenStudy (anonymous):

Divide Both Sides By -2.

OpenStudy (anonymous):

so its -50?

OpenStudy (anonymous):

Positive. Two Negatives Equal A Positive.

OpenStudy (anonymous):

ok so they sell 50 each?

OpenStudy (anonymous):

No. Yu only found C. So we have to go back and plug in the answer to find the amount of ham sandwiches. So take the first equation. 120=H+C(50)

OpenStudy (anonymous):

Subtract 50 from both sides... 120-50=h+50-50 And you get H=70.

OpenStudy (anonymous):

so Ham Sandwiches sold 70 and the chicken sold 100?

OpenStudy (anonymous):

No No No. 50 chicken.

OpenStudy (anonymous):

Remember. It has to equal 120. 120=70+50

OpenStudy (anonymous):

ok i get it now. ill review it again later. how about the other one?

OpenStudy (anonymous):

That one is seriously throwing me off. Can you post that question again?

OpenStudy (anonymous):

its obvious i didnt copy it correctly.

OpenStudy (anonymous):

so nvm that thank you for your help

OpenStudy (anonymous):

Welcome! Good Luck!

OpenStudy (anonymous):

btw what is let x = in no 1?

OpenStudy (anonymous):

What Do yu Mean?

OpenStudy (anonymous):

where did you get h+50-50

OpenStudy (anonymous):

Ohh That Was Me Subtracting %0 From Both Sides. 120=h+50 -50 -50 Is that better?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No Problem.